What is the sum of the first 10 terms of the sequence defined by an=3n-3?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

What is the sum of the first 10 terms of the sequence defined by an=3n-3?
-------
a(1) = 3-3 = 0
a(2) = 3*2-3 = 3
a(3) = 3*3-3 = 6
-----
a(10) = 3*10-3 = 27
----------------------
Sum of 1st 10::
S(10) = (10/2)(0+27) = 5*27 = 135
------------------------------------
Cheers,
Stan H.
------------------

Answer by richwmiller(17219)

an=3n-3=3*(n-1)
a1=3-3=0
a2=3*2-3=3
a3=3*3-3=6
a4=3*4-3=9
a5=3*5-3=12
a6=3*6-3=15
a7=3*7-3=18
a8=3*8-3=21
a9=3*9-3=24
a10=3*10-3=27
0,3,6,9,12,15,18,21,24,27=135
S = ½(2a + (n-1)d)n
S = ½(2*0 + (9)3)10=135